OpenStudy (anonymous):

find y" of xy +tan exponent of-1(xy) =4

OpenStudy (dumbcow):

is that find 2nd derivative?

OpenStudy (anonymous):

yes could you help me?

OpenStudy (dumbcow):

well you need to find 1st derivative first using implicit differentiation \[d/dx \tan^{-1} u = \frac{1}{1+u^{2}}*du\] where u = xy in this case (xy)' = y + xy' \[\rightarrow y+xy' +\frac{y+xy'}{1+x^{2}y^{2}} = 0\] \[(y+xy')(1+x^{2}y^{2}) + (y+xy') = 0\] \[\rightarrow y+xy' = 0\] \[y' =- \frac{y}{x}\] Now differentiate again using quotient rule \[y'' = \frac{-xy'+y}{x^{2}}\] substitute in for y' \[y'' = \frac{2y}{x^{2}}\]